### v2

# KANELÉ: Kolmogorov–Arnold Networks for Efficient LUT-based Evaluation

Abstract.
Low-latency, resource-efficient neural network inference on FPGAs is essential for applications demanding real-time capability and low power. Lookup table (LUT)-based neural networks are a common solution, combining strong representational power with efficient FPGA implementation. In this work, we introduce KANELÉ, a framework that exploits the unique properties of Kolmogorov–Arnold Networks (KANs) for FPGA deployment. Unlike traditional multilayer perceptrons (MLPs), KANs employ learnable one-dimensional splines with fixed domains as edge activations, a structure naturally suited to discretization and efficient LUT mapping. We present the first systematic design flow for implementing KANs on FPGAs, co-optimizing training with quantization and pruning to enable compact, high-throughput, and low-latency KAN architectures. Our results demonstrate up to a 2700x speedup and orders of magnitude resource savings compared to prior KAN-on-FPGA approaches. Moreover, KANELÉ matches or surpasses other LUT-based architectures on widely used benchmarks, particularly for tasks involving symbolic or physical formulas, while balancing resource usage across FPGA hardware. Finally, we showcase the versatility of the framework by extending it to real-time, power-efficient control systems.
Kolmogorov–Arnold Networks (KANs), FPGAs, Lookup tables (LUTs), Neural networks, Quantization, Pruning, Hardware–software codesign
journalyear: 2026copyright: ccconference: Proceedings of the 2026 ACM/SIGDA International Symposium on Field Programmable Gate Arrays; February 22–24, 2026; Seaside, CA, USAbooktitle: Proceedings of the 2026 ACM/SIGDA International Symposium on Field Programmable Gate Arrays (FPGA ’26), February 22–24, 2026, Seaside, CA, USAdoi: 10.1145/3748173.3779202isbn: 979-8-4007-2079-6/2026/02ccs: Computing methodologies Machine learning algorithmsccs: Hardware
Figure 1. From the Kolmogorov-Arnold Representation Theorem to efficient KAN FPGA inference.
1. Introduction
Lookup table (LUT) based neural networks have become a central paradigm for efficient FPGA inference, with designs such as NeuralLUT-Assemble (Andronic and Constantinides, 2025), TreeLUT (Khataei and Bazargan, 2025), DWN (Bacellar et al., 2025), and others (Andronic and Constantinides, 2023; Weng et al., 2025; Chang et al., 2024; Sun et al., 2025) demonstrating dramatic gains in area, latency, and power efficiency. These approaches highlight the advantages of rethinking neural networks around LUT primitives, though they remain largely confined to supervised learning and task-specific architectures.
In this work, we demonstrate that Kolmogorov–Arnold Networks (KANs) offer a principled foundation for LUT-based design. Inspired by the Kolmogorov–Arnold representation theorem, KANs replace the fixed activations of Multilayer Perceptrons (MLPs) with learnable edge functions and the matrix multiplication in MLPs with summation at nodes (Fig. 1). This activation-centric formulation aligns naturally with LUTs: each learnable spline defined on a fixed domain can be quantized, pruned, and directly mapped to LUTs. In the literature, although KANs have been shown to outperform MLPs in settings such as PDE solving and scientific computing (Liu et al., 2025, 2024), their practical deployment has been hindered by slow inference and costly hardware realizations (Tran et al., 2024; Hou et al., 2025). The only prior FPGA implementation concluded KANs were impractical, due to expensive spline evaluations and high resource usage (Tran et al., 2024).
This paper shows that by re-formulating KAN inference entirely in terms of LUTs, KANs are not only feasible but highly efficient in FPGA settings. Thus, our contributions are fourfold:
  1. (1)
    FPGA-tailored KAN Architecture: We present KANELÉ, named after the French pastry known for its compact form and rich structure (Marmiton, n.d.). At its core, the framework co-optimizes quantization, pruning, and mapping of KAN functions onto learned LUTs and additions, thereby minimizing memory and logic overhead. From a KAN research perspective, KANELÉ is the first FPGA-tailored formulation, eliminating BRAM/DSP usage, reducing latency by up to $2700\times$, and cutting resource usage by over $4000\times$ compared to prior designs (Tran et al., 2024).
  2. (2)
[…]
  3. (3)
    Open-source Framework: We provide an automated software–hardware co-design flow that compiles KANs into optimized FPGA implementations within seconds, supporting reproducible studies across domains such as biology, physics, vision, signal processing, and tabular ML. Code is available at: https://github.com/Duchstf/KANELE
  4. (4)
    Control Systems: we extend KANELÉ beyond supervised learning to continuous control, showing on the HalfCheetah benchmark from OpenAI Gym (Todorov et al., 2012) that a quantized KAN policy with $\sim 5\times$ fewer parameters than an MLP baseline policy achieves higher rewards, underscoring its suitability for resource-constrained, real-time control systems.
2. Background & Related Works
This section reviews Kolmogorov–Arnold Networks (KANs) and prior work on LUT-based neural network inference.
2.1. Kolmogorov–Arnold Networks
Kolmogorov–Arnold Networks (KANs) replace the fixed activation functions and matrix multiplications of MLPs with learnable spline-based functions on network edges (Liu et al., 2025). This activation-centric formulation improves expressiveness and interpretability, often achieving comparable accuracy with fewer parameters and operations. Since their introduction, KANs have inspired extensive follow-up work, including theoretical analyses (Yu et al., 2024; Wang et al., 2025a), architectural extensions (e.g., convolutional (Bodner et al., 2025; Drokin, 2024), temporal (Genet and Inzirillo, 2025), and Fourier-based variants (Zhang et al., 2025; Imran and Ishmam, 2024)), and applications across scientific modeling and data-driven tasks (Wang et al., 2025b; Cruz et al., 2025; Li et al., 2024; Han et al., 2025).
Despite this rapid progress, recent surveys identify computational efficiency and hardware implementation as key open challenges (Ji et al., 2025; Somvanshi et al., 2025; Hou et al., 2025). To date, efficient hardware realization remains largely unexplored, with one early attempt concluding that a direct FPGA implementation incurs prohibitive resource and latency costs compared to MLPs (Tran et al., 2024). Our work directly challenges this conclusion by demonstrating that the activation-centric design of KANs is, in fact, exceptionally well-suited for hardware acceleration through a LUT-based paradigm.
2.2. LUT-based Neural Networks
LUT-based neural networks aim to replace arithmetic-heavy MAC operations with precomputed function evaluations stored in LUTs, exploiting the abundant and low-latency logic resources of FPGAs. Pioneering frameworks like LUTNet (Wang et al., 2019) and LogicNets (Umuroglu et al., 2020) first demonstrated the replacement of arithmetic with direct LUT mappings. Subsequent works generalized this concept to approximate more complex functions (Andronic and Constantinides, 2023, 2024) and improved scalability using additive or modular ensembles (Lou et al., 2024; Weng et al., 2025; Andronic and Constantinides, 2025). This design philosophy has also been used to efficiently implement other machine learning models, such as gradient-boosted decision trees (Khataei and Bazargan, 2025). Another related approach is the family of Weightless Neural Networks (WNNs), which stores learned patterns directly in LUTs (Aleksander et al., 1984; Miranda et al., 2022; Susskind et al., 2023, 2022; Bacellar et al., 2025), though often at the cost of representational power.
Conceptually, KANs are close to PolyLUT (Andronic and Constantinides, 2023), PolyLUT-Add (Lou et al., 2024), and DWNs (Bacellar et al., 2025) but possess distinct structural properties. While PolyLUT tabulates multivariate polynomials, which theoretically allows the native representation of arbitrary products $p(\mathbf{x})=\prod_{i}x_{i}$, this approach suffers from exponential LUT growth relative to input dimension. In contrast, KANs decompose functions into sums of tabulated univariate splines. Although this formulation relies on layer composition to approximate the multiplicative terms inherent to PolyLUT, it yields linear scaling with input dimension and an additive structure that is naturally amenable to pruning. While this formulation doesn’t explicitly represent pure multiplicative terms, in practice, compositions of low-order ($\leq$3) splines approximate such interactions effectively. Moreover, DWN’s full binarization of inputs and LUTs hinders generalization beyond classification, while KANELÉ supports higher-precision arithmetic for tasks such as autoencoding and continuous control. Finally, DWN’s finite-difference differentiability may further constrain optimization flexibility compared to KANELÉ gradient descent.
3. KAN Architecture and Quantization-Aware Training and Pruning
We design the KANELÉ framework for KAN FPGA deployment using quantization-aware training and pruning, enabling efficient hardware translation while preserving consistency between training and inference.
3.1. KAN Architecture with Learnable Activation Functions
Before introducing quantization and pruning, we first outline the core architecture. Unlike MLPs, KANs replace fixed nonlinearities with learnable activation functions, each modeled as a linear combination of B-spline basis functions with trainable coefficients. B-splines are piecewise polynomials defined on a grid, providing smoothness, locality, and efficient nonlinear parameterization. More generally, activations can be expressed in other orthogonal bases, such as Fourier series (Li et al., 2025; Zhang et al., 2025; Imran and Ishmam, 2024).
A KAN layer with $d_{\mathrm{in}}$ inputs and $d_{\mathrm{out}}$ outputs is represented as a matrix of 1D learnable functions
(1) $\Phi=\{\phi_{q,p}\},\quad p=1,\dots,d_{\mathrm{in}},\quad q=1,\dots,d_{\mathrm{out}},$
where each $\phi_{q,p}$ is trainable (Fig. 2).
Figure 2. A KAN activation $\phi(x)$ represented as a linear combination of B-spline basis functions $B_{i}(x)$ on a grid over $[a,b]$: $\phi(x)=\sum_{i}c_{i}B_{i}(x)$. Trainable coefficients $c_{i}$ control the overall function shape.
For improved convergence, each $\phi_{q,p}$ combines a base activation $\phi(\cdot)$ with B-splines $\{B_{p,k}(\cdot)\}$:
(2) $\phi_{q,p}(x_{p})=w^{\mathrm{base}}_{q,p}\,\phi(x_{p})+\sum_{k=1}^{G+S}w^{\mathrm{spline}}_{q,p,k}\,B_{p,k}(x_{p}),$
where $w^{\mathrm{base}}_{q,p}$ and $w^{\mathrm{spline}}_{q,p,k}$ are trainable. Splines are defined on a grid of size $G$ and order $S$ within a fixed domain $[a,b]$.
Given $x_{l}\in\mathbb{R}^{d_{\mathrm{in}}}$, the output is
(3) $(x_{l+1})_{j}=\sum_{i=1}^{d_{l}}\phi_{j,i}(x_{l,i}),\quad j=1,\dots,d_{l+1},$
or in compact form
(4) $x_{l+1}=\Phi_{l}(x_{l}),$
with $\Phi_{l}$ the function matrix of layer $l$. A $L$-layer KAN is thus
(5) $\mathrm{KAN}(x)=\Phi_{L-1}\circ\Phi_{L-2}\circ\cdots\circ\Phi_{0}(x).$
As illustrated in Fig. 1, KANs extend MLPs by learning activations directly, offering greater representational flexibility while preserving a structured, layer-wise graph.
3.2. Quantization-Aware Training
For efficient FPGA deployment, we adopt quantization-aware training (QAT) via AMD’s Brevitas library (Franco et al., 2025). Quantizers are placed at the network input and after each KAN layer, ensuring that training adapts to the required hardware precision.
For a layer $l$ with output $\mathbf{x}_{l+1}\in\mathbb{R}^{d_{l+1}}$, the quantized output is
(6) $\mathbf{x}_{l+1,q}=q_{l}(\mathbf{x}_{l+1}),$
where $q_{l}(\cdot)$ is the layer quantizer. Similarly, an input quantizer $q_{I}(\cdot)$ is applied to $\mathbf{x}_{0}$.
The layer output quantizer performs $n_{l}$-bit uniform quantization:
(7) $\mathbf{x}_{l+1,q}=s_{l}\cdot\text{Quantize}[n_{l}]\!\left(\frac{\operatorname{clip}(\mathbf{x}_{l+1},a,b)}{s_{l}}\right),$
where $s_{l}$ is a learnable scale (fixed at inference), and $[a,b]$ is the shared quantization domain (Fig. 3).
Figure 3. Layer-wise uniform quantization. Here, 2-bit inputs $q_{l-1}(x_{l})$ are mapped to 3-bit outputs $q(x_{l+1})$ over the fixed range $[a,b]$. Orange markers indicate quantization levels; the dotted curve is the underlying continuous mapping.
The input quantizer incorporates both scale $s_{I}$ and bias $b_{I}$ to handle asymmetric distributions:
(8) $\mathbf{x}_{0,q}=s_{I}\cdot\text{Quantize}[n_{I}]\!\left(\frac{\operatorname{clip}(\mathbf{x}_{0},a,b)}{s_{I}}+b_{I}\right).$
During RTL generation, $s_{I}$ and $b_{I}$ are fixed for deterministic behavior.
In practice, input preprocessing is realized by a batch normalization (zero mean, unit variance) followed by a ScalarBiasScale block introducing $b_{I}$ and $s_{I}$. At inference, BN statistics are folded into these constants, yielding an affine shift–scale, clipping, and quantization. This design preserves LUT-based compatibility while avoiding the overhead of full batch normalization.
During training, quantizer gradients are approximated using the straight-through estimator (STE):
(9) $\frac{\partial q(x)}{\partial x}\approx 1,$
which allows gradient flow through quantized operations without modification.
3.3. Pruning via Norm-Based Selection
To reduce resource usage, we prune spline connections by evaluating their contribution over the input domain. In contrast to conventional LUT-based neural networks—which rely on sequential LUT indexing, making every LUT entangled with the next and thus nearly impossible to prune without breaking the model—KANELÉ exploits the inherently additive structure of KANs, where each LUT contributes independently to a summation. This independence makes pruning both mathematically natural and directly compatible with FPGA hardware. The original KAN paper emphasizes efficient pruning as a key advantage of KAN’s edge-centric architecture, and we extend this insight to demonstrate a distinct advantage over node-based LUT networks for efficient hardware translation (Liu et al., 2025).
For each pair $(i,j)$ of input and output neurons, we compute the activation of the spline component:
(10) $f_{p\to q}(x)=\sum_{k=1}^{G+S}w^{\mathrm{spline}}_{q,p,k}\,B_{p,k}(x).$
Its importance is measured via the $\ell_{2}$ norm across a sampled input grid $\mathcal{X}$ consistent with its quantization level:
(11) $\|f_{p\to q}\|_{2}=\left(\sum_{x\in\mathcal{X}}\left|f_{p\to q}(x)\right|^{2}\right)^{1/2}.$
A structured pruning mask is then applied:
(12) $m_{q,p}=\begin{cases}1,&\|f_{p\to q}\|_{2}>\tau(t),\\ 0,&\text{otherwise},\end{cases}$
where $\tau(t)$ is a pruning threshold that changes as a function of epochs ($t$) with
$\tau(t)=T\text{exp}\left(-\ln 20\cdot\dfrac{\max(t,t_{0})}{t_{f}-t_{0}}\right).$
This pruning threshold corresponds to an exponential warmup, where pruning starts on epoch $t_{0}$ and increases exponentially, hitting $95\%$ of the full pruning threshold $T$ on target epoch $t_{f}$. This allows us to control pruning dynamics to avoid interference with proper training. Backward pruning is additionally applied if the corresponding output neuron has no active connections in the subsequent layer, ensuring consistent sparsity propagation.
3.4. KAN Hyperparameters
To summarize, the training and deployment of KANs involves hyperparameters which can be broken up into three main classes: spline representation hyperparameters, hardware architecture hyperparameters, and pruning hyperparameters. The descriptions and impact of each hyperparameter are detailed in Table 1. The joint optimization of these parameters provides a flexible design space that balances learning capacity with hardware efficiency in FPGA deployments.
Table 1. Summary of KAN training parameters. The first group influences the accuracy through spline representation, the second group encodes the hardware architecture, and the third group determines pruning policy which affects hardware architecture.
Symbol Description Impact
$G$ Grid size (number of intervals) Controls spline resolution; accuracy only
$[a,b]$ Grid range (spline domain) Defines support of basis functions; accuracy only
$S$ Spline order Smoothness and flexibility; accuracy only
$d_{l}$ Layer dimensions Affects model capacity and resource usage
$n_{l}$ Layer bitwidth (QAT precision) Direct trade-off between accuracy and resource cost
$T$ Pruning threshold Governs sparsity and LUT reduction and accuracy tradeoff
$t_{0}$ Warmup start epoch Determines when pruning warmup starts
$t_{f}$ Warmup target epoch Affects pruning warmup rate
4. LUT-Based KAN Architecture
This section outlines our end-to-end mapping of trained KANs to synthesizable VHDL RTL and associated pipelining strategies. Our end-to-end toolflow currently supports the basic KAN architecture using B-splines, as in the original paper (Liu et al., 2025). Extensions to other bases or architectures such as convolutions or transformers are feasible. The framework is designed for usability—anyone familiar with training MLPs can readily train and deploy KANELÉ.
4.1. Toolflow
A high-level overview of the toolflow stages is shown in Figure 4. This push-button workflow removes manual RTL work: starting from a trained PyTorch checkpoint, it deterministically emits RTL, memory images, and build/simulation scripts, enabling rapid deployment of arbitrary KAN topologies to FPGA and bitstream generation via Vivado.
Figure 4. Visualization of the KAN to FPGA implementation toolflow.
4.1.1. Training
The training process begins by specifying the KAN hyperparameters (see Section 3.4) together with the dataset to be used. The user may freely select the optimizer and learning-rate schedule best suited to their task. In our implementation, we adopt the PyTorch AdamW optimizer (Loshchilov and Hutter, 2019) as the default choice due to its robustness in handling weight decay. The model is then trained with the QAT and pruning mechanism described in Section 3.2 and Section 3.3, respectively, to reduce resource usage and latency, while preserving accuracy. This step produces compact learned activations for KAN that can be efficiently mapped to LUT representations for FPGA deployment.
4.1.2. KAN to Logical-LUTs Conversion
Following (Andronic and Constantinides, 2025), we denote lookup tables extracted from the network as Logical-LUTs (L-LUTs), and FPGA fabric resources as Physical-LUTs (P-LUTs). From a trained, pruned, and quantized PyTorch KAN, each surviving connection is translated into an L-LUT. For every active edge, the input state space is enumerated and the KAN layer’s pre-activation response is evaluated and quantized. This produces per-connection truth tables stored as compact JSON files, yielding a deterministic, bit-accurate mapping of the model into integer-valued L-LUTs. The representation preserves quantization and sparsity, enabling efficient FPGA deployment.
4.1.3. RTL File Generation
From the L-LUT graph, we generate a complete RTL design for FPGA deployment. The tool emits VHDL sources for the KAN core, per-layer packages, LUT entities, and memory initialization files encoding the truth tables. A configuration package specifies bit widths, signal types, and accumulator sizes. Each L-LUT is instantiated as a memory-mapped component, organized into layers, with balanced adder trees for output accumulation. Pipeline registers are inserted between layers to improve clocking and shorten critical paths. The result is a self-contained firmware bundle that includes simulation testbenches, initialization vectors, and Vivado build scripts—supporting functional simulation, latency evaluation, and FPGA synthesis.
4.1.4. Synthesis and Place & Route
We synthesize the generated RTL with Vivado 2024.1, targeting the xcvu9p-flgb2104-2-i FPGA for benchmarking against LUT-based networks, and the xczu7ev-ffvc1156-2-e FPGA for comparison with prior KAN works. To ensure fairness and consistency with prior works, we use settings that isolate core delay and area, specifically Vivado’s Flow_PerfOptimized_high mode with Out-of-Context synthesis, allowing each module to be compiled independently. While the maximum clock is ultimately limited by the FPGA’s global clock, LUT-centric designs such as KANELÉ  typically sustain high frequencies, making the computational core unlikely to be the critical path in larger systems. The target clock period is therefore chosen relative to network size, following prior work.
4.2. Pipelining Strategies
Efficient pipelining is crucial for achieving high FPGA clock frequencies while maintaining low latency across KAN layers. We introduce pipelining at two levels: (i) within adder trees that accumulate outputs of multiple Logical-LUTs (L-LUTs) per channel, and (ii) between consecutive network layers.
Adder Tree Pipelining.
Each neuron computes a weighted sum of active inputs via a reduction tree over L-LUT outputs. A naïve single-stage sum creates a long combinational path, thereby limiting frequency. Instead, we implement a balanced, pipelined adder tree with registers after each stage. At each stage up to $n_{\mathrm{add}}$ inputs are combined, reducing fan-in and distributing additions over multiple cycles. The depth is
$\text{depth}_{\ell}=\left\lceil\log_{n_{\mathrm{add}}}(N_{\ell})\right\rceil,$
as shown in Figure 5 for $n_{\mathrm{add}}=2$. At the end of the adder tree, quantization and saturation of the sum are performed to make the output consistent with the subsequent layer’s input. This is taken into account during training, preventing any degradation in accuracy.
Figure 5. Balanced, pipelined adder tree for computing one neuron activation with $n_{add}=2$.
Inter-Layer Pipelining.
Pipeline registers are also inserted between layers, capturing saturated outputs before feed-forward. This isolates LUT evaluation, summation, and activation in time, minimizing critical paths and balancing latency. Register insertion is automated in RTL generation, ensuring deep pipelining for arbitrary KAN topologies.
Limitations.
As a LUT-based model, KAN inherits known limitations. LUT size scales exponentially with input bitwidth, though only linearly with fan-in (Andronic and Constantinides, 2023, 2024). For high-dimensional inputs, preserving KAN’s structure requires long adder chains, reducing throughput and increasing resource usage. Adder trees sustain high clock frequencies by adding pipeline stages, at the cost of extra clock cycles. Thus, image tasks like MNIST require aggressive pruning to remain resource-feasible, with some accuracy loss.
5. Experimental Results
5.1. Benchmarks
We evaluate KAN across three domains of datasets: (i) widely adopted benchmark datasets from the LUT-based neural network literature, (ii) synthetic and tabular datasets previously used in KAN FPGA benchmarking (Tran et al., 2024), and (iii) MLPerf Tiny datasets, which provide real-world tasks with more complex modalities and objectives. Together, these domains form a diverse testbed to study KAN across different tasks and levels of complexity. Below we briefly describe each dataset.
5.1.1. LUT-based Neural Network Benchmarks
  • MNIST: A large-scale handwritten digit recognition dataset containing 60,000 training and 10,000 test images of size $28\times 28$, labeled across 10 classes (digits 0–9) (Deng, 2012).
  • JSC OpenML: A tabular dataset (CERN Collaboration, 2025) from the JSC suite, consisting of 16 jet substructure input features and a 5-class jet classification task. It has been widely used in comparisons of LUT-based networks. This version contains 830,000 instances and is known to exhibit easier convergence, possibly due to improved data curation (Andronic and Constantinides, 2025).
  • JSC CERNBox: Another dataset (OpenML Contributors and LHC Jets HLF Curators, 2020) from the JSC benchmark, involving the same jet tagging task. It comprises 986,806 instances and is generally considered more challenging, as models trained on it tend to achieve lower accuracies due to the increased dataset complexity.
5.1.2. KAN FPGA Benchmarks
  • Moons: A synthetic two-class dataset (Pedregosa et al., 2011) commonly used for testing nonlinear decision boundaries. Each point lies in one of two interleaving half-moon shapes with added Gaussian noise.
  • Wine: A dataset from the University of California, Irvine (UCI) Machine Learning Repository (Aeberhard and Forina, 1992), containing 13 physicochemical attributes of wine samples classified into 3 quality categories.
  • Dry Bean: Another UCI dataset (15) with 16 numerical features representing bean shape and texture, used for classifying 7 different bean varieties .
5.1.3. MLPerf Tiny Benchmarks
  • ToyADMOS: An audio anomaly detection dataset featuring sound files from both normally functioning and defective toy cars. An autoencoder is trained on sliding windows of the downsampled mel spectrogram (input size 64). For classification, the mean reconstruction loss across all sliding window spectrograms of an audio file is computed, and a fixed threshold is applied to label the sample as an anomaly. Notably, this benchmark presents significantly more complex inputs compared to the LUT-NN and KAN-FPGA benchmarks while also utilizing a non-classification objective (reconstruction loss). (Banbury et al., 2021).
These datasets span a spectrum of complexity, from low-dimensional toy datasets (Moons) to high-dimensional real-world data (JSC, MNIST, ToyADMOS), ensuring that both the representational capacity and hardware efficiency of KANELÉ are thoroughly evaluated.
Table 2. Accuracy comparison of MLP Floating Point and KAN Floating Point and Quantized models on benchmark datasets. All models have the same dimensions listed. The floating point versions only use the layer size ($d_{l}$) parameter.
Datasets $G$ $[a,b]$ $S$ $d_{l}$ $n_{l}$ $T$ Accuracy (%)/AUC
MLP FP KAN FP KAN Quantized & Pruned
KAN FPGA Benchmarks (Tran et al., 2024)
Moons 6 [-8, 8] 3 [2, 2, 1] [6, 5, 8] 0. 87.2 97.7 97.4
Wine 6 [-8, 8] 3 [13, 4, 3] [6, 7, 8] 0. 96.3 98.1 98.2
Dry Bean 6 [-8, 8] 3 [16, 2, 7] [6, 6, 8] 0. 90.9 92.2 92.1
LUT-based neural network benchmarks
MNIST 30 [-8,8] 3 [784, 62, 10] [1, 6, 6] 1. 96.7 97.9 96.3
JSC CERNBox 30 [-2, 2] 10 [16, 12, 5] [8, 8, 6] 0.14 73.0 75.1 75.1
JSC OpenML 40 [-2, 2] 10 [16, 8, 5] [6, 7, 6] 0.9 76.5 76.5 76.0
MLPerf Tiny Benchmark
ToyADMOS 30 [-2, 2] 10 [64, 16, 8, 16, 64] [7, 8, 8, 7, 8] 0.9 0.80 0.83 0.83
5.2. Training Parameters
KAN hyperparameters are straightforward to manage. The spline-related parameters ($G,[a,b],S$) only affect accuracy and can often be set to robust defaults since they do not impact hardware resources.
In practice, balancing model performance with hardware efficiency centers on tuning three key parameters: the layer dimensions ($d_{l}$), bitwidth ($n_{l}$), and pruning threshold ($T$). These directly control the model’s capacity, numerical precision, and sparsity. As demonstrated in Table 2, this allows the quantized and pruned KANs to achieve competitive accuracy—even outperforming floating-point versions on datasets like Wine (98.2%)—while being optimized for an efficient FPGA implementation.
5.3. Comparison with LUT-NN Architectures
We benchmarked KANELÉ against state-of-the-art LUT-based architectures on three datasets: JSC CERNBox, JSC OpenML, and MNIST (Table 3). It should be noted that, as detailed in Table 2, we assume the same input bitwidth compared to prior works, which is not the case for DWN (Bacellar et al., 2025) that uses a thermometer encoding to assign distinct floating-point thresholds to each feature, leading to potentially large overhead.
JSC CERNBox
[…]
JSC OpenML
[…]
MNIST
[…]
Table 3. Evaluation of KANELÉ against state-of-the-art ultra-low-latency, resource-efficient LUT-based network architectures. Results are reported after performing out-of-context synthesis and place-and-route. Input bit-widths are consistent with those used in prior works for fair comparison.
[…]
JSC CERNBox KANELÉ 75.1 5034 1917 0 0 870 8.1 $4.1\times 10^{4}$
NeuraLUT-Assemble (Andronic and Constantinides, 2025) 75.0 8539 1332 0 0 352 5.7 $4.87\times 10^{4}$
AmigoLUT-NeuraLUT (Weng et al., 2025) 74.4 42742 4717 0 0 520 9.6 $4.10\times 10^{5}$
PolyLUT-Add (Lou et al., 2024) 75.0 36484 1209 0 0 315 16 $5.84\times 10^{5}$
NeuraLUT (Andronic and Constantinides, 2024) 75.1 92357 4885 0 0 368 14 $1.29\times 10^{6}$
PolyLUT (Andronic and Constantinides, 2023) 75.0 246071 12384 0 0 203 25 $6.15\times 10^{6}$
LogicNets (Umuroglu et al., 2020) 72.0 37931 810 0 0 427 13 $4.93\times 10^{5}$
JSC OpenML KANELÉ 76.0 1232 900 0 0 987 7.1 $8.7\times 10^{3}$
NeuraLUT-Assemble (Andronic and Constantinides, 2025) 76.0 1780 540 0 0 941 2.1 $3.92\times 10^{3}$
TreeLUT (Khataei and Bazargan, 2025) 75.6 2234 347 0 0 735 2.7 $6.03\times 10^{3}$
DWN (Bacellar et al., 2025) 76.3 4972 3305 0 0 827 7.3 $3.6\times 10^{4}$
da4ml (Sun et al., 2025) 76.9 12250 1502 0 0 212 18.9 $2.3\times 10^{5}$
hls4ml (Fahim et al.) (Fahim et al., 2021) 76.2 63251 4394 38 0 200 45 $2.85\times 10^{6}$
MNIST KANELÉ 96.3 3809 4133 0 0 864 9.3 $3.5\times 10^{4}$
NeuraLUT-Assemble (Andronic and Constantinides, 2025) 97.9 5070 725 0 0 863 2.1 $1.06\times 10^{4}$
TreeLUT (Khataei and Bazargan, 2025) 96.6 4478 597 0 0 791 2.5 $1.12\times 10^{4}$
DWN (Bacellar et al., 2025) 97.8 2092 1757 0 0 873 9.2 $1.92\times 10^{4}$
PolyLUT-Add (Lou et al., 2024) 96.0 14810 2609 0 0 625 10 $1.48\times 10^{5}$
AmigoLUT-NeuraLUT (Weng et al., 2025) 95.5 16081 13292 0 0 925 7.6 $1.22\times 10^{5}$
NeuraLUT (Andronic and Constantinides, 2024) 96.0 54798 3757 0 0 431 12 $6.58\times 10^{5}$
PolyLUT (Andronic and Constantinides, 2023) 97.5 75131 4668 0 0 353 17 $1.38\times 10^{6}$
FINN (Umuroglu et al., 2017) 96.0 91131 0 5 200 310 $2.82\times 10^{7}$
hls4ml (Ngadiuba et al.) (Ngadiuba et al., 2020) 95.0 260092 165513 0 345 200 190 $4.94\times 10^{7}$
Table 4. FPGA resource utilization and latency of KAN models on Moons, Wine, and Dry Bean benchmarks as used in (Tran et al., 2024)
[…]
Moons KANELÉ 97 1736 0 0 57 67 5 2.9 $1.9\times 10^{2}$
KAN (Tran et al) (Tran et al., 2024) 97 - 10 120 8622 17877 128 1280 $2.3\times 10^{7}$
ChebyUnit (Zhonglongyou et al., 2026) 100 - 10 40 12150 9888 13 130 $1.3\times 10^{6}$
Wine KANELÉ 98 983 0 0 686 534 6 6.1 $8.8\times 10^{3}$
KAN (Tran et al) (Tran et al., 2024) 97 - 132 950 74741 146843 688 6880 $1.0\times 10^{9}$
ChebyUnit (Zhonglongyou et al., 2026) 95 - 132 324 22104 30154 13 130 $3.9\times 10^{6}$
Dry Bean KANELÉ 92 842 0 0 471 402 6 7.1 $3.3\times 10^{3}$
KAN (Tran et al) (Tran et al., 2024) 92 - 781 9111 734544 1677558 1896 18960 $3.2\times 10^{10}$
ChebyUnit (Zhonglongyou et al., 2026) 92 - 781 256 25198 27359 13 130 $3.6\times 10^{6}$
Table 5. Comparison of FPGA resource utilization, latency, and power consumption for the anomaly detection task on the ToyADMOS time series dataset in the MLPerf Tiny Benchmark, evaluated on the xc7a100t-1csg324 FPGA (Banbury et al., 2021).
Dataset Model AUC BRAM (36kb) DSP $F$ LUT LUTRAM II (clocks) Throughput (inf/s) Latency ($\mu s$) Energy/inf. ($\mu J$)
ToyADMOS KANELÉ 0.83 0 0 17,643 29,981 0 1 228 M 0.07 0.01
hls4ml (MLPerf Tiny v0.7) (Borras et al., 2022) 0.83 22.5 207 61,639 51,429 5,780 144 694 k 45 98.4
5.4. Comparison with Prior KAN-FPGA Literature
We benchmarked KANELÉ against the KAN FPGA implementation by Tran et al. (Tran et al., 2024) on the Moons, Wine, and Dry Bean datasets. The results, detailed in Table 4, demonstrate that KANELÉ offers a dramatic improvement in hardware efficiency and performance while maintaining or exceeding the accuracy of the previous work.
Through its LUT-based approach, our implementation completely eliminates the need for BRAM and DSP blocks, which are heavily utilized in the implementation by Tran et al. This leads to a massive reduction in the overall hardware footprint. For instance, on the Dry Bean dataset, KANELÉ uses only 402 LUTs and 471 FFs, whereas the previous work consumes over 1.6 million LUTs and 734,000 FFs—a reduction of more than 4000x in LUTs.
[…]
Figure 6. Ablation study of KANELÉ on the JSC OpenML benchmark demonstrating trade-offs between accuracy, pruning, and resource usage. (a) Accuracy improves steadily as hardware resources scale up with LUT and FF usage growing at roughly the same rate. (b) LUT/$F$ usage scales roughly linearly with the number of unpruned edges. (c) LUT/$F$ usage scales linearly with hidden layer width, confirming a direct mapping from learned activation functions (edges) to resources. (d) Decreasing activation bitwidths reduces LUT resource usage exponentially, with diminishing returns observed below 6 bits.
5.5. Comparison with hls4ml MLPerf Tiny
To understand the performance of KANELÉ on more complex datasets, we compare the framework to hls4ml (Borras et al., 2022) on the ToyADMOS dataset, part of the MLPerf Tiny benchmark suite. Other LUT-based neural networks have not attempted this benchmark, possibly due to its high complexity and/or non-classification-based training scheme. The results in Table 5 indicate that KANELÉ achieves substantially better performance than prior approaches in terms of both resource efficiency, latency, and power. These improvements highlight the potential of KANELÉ for future studies involving more complex datasets as well as tasks beyond classification. Specifically, KANELÉ eliminates the need for BRAM, LUTRAM, and DSPs, while reducing LUT usage by 41.7% and FF usage by 71.4% relative to hls4ml. In terms of performance, KANELÉ delivers $330\times$ higher throughput, $643\times$ lower latency and a $9,840\times$ reduction in energy per inference. These results establish KANELÉ as a highly efficient alternative to existing FPGA neural implementations, with strong potential for scaling to more complex datasets and tasks beyond classification.
5.6. Ablation Study
We perform an ablation study for KANELÉ using the JSC OpenML dataset, shown in Figure 6. This analysis isolates the effect of four key design factors—accuracy, pruning, hidden layer width, and model bitwidth—on FPGA resource utilization (LUTs and FFs). It is seen that pruning and quantization provide the most effective levers for controlling hardware footprint, while hidden layer width and accuracy tuning allow fine-grained trade-offs between performance and efficiency. Another important conclusion is that the size of a KAN network can holistically be thought of in terms of its number of edges: this is proportional to the number of LUTs and FFs used, as seen in Figures  6(b) and  6(c). Together, these insights guide principled design choices for deploying KANELÉ under tight FPGA resource budgets.
5.7. Extension to Real-time Control Systems
To demonstrate that the KANELÉ paradigm extends well beyond the traditional supervised learning tasks typically studied in the LUT-based neural network community, we apply our framework to a reinforcement learning benchmark. Specifically, we evaluate on the HalfCheetah environment, a classic continuous control task in reinforcement learning (RL), most commonly accessed through the MuJoCo physics simulator via OpenAI Gym (now Gymnasium) (Todorov et al., 2012). The objective is to learn a policy that enables a simulated two-legged agent to run as fast and as stably as possible. This environment is widely used as a standard benchmark to compare RL algorithms and function approximators. While HalfCheetah is a simulated benchmark, it captures key aspects of many practical robot control tasks, as they all rely on the same underlying design and physics principles.
This extension is motivated by prior work such as (Kich et al., 2024), where the authors showed that KAN actor networks with significantly fewer trainable parameters achieved higher rewards compared to much larger MLPs when trained with the Proximal Policy Optimization (PPO) algorithm. PPO is one of the most widely used RL algorithms in practice, particularly in robotics, games, and continuous control tasks.
5.7.1. Setup
Since the primary focus of this paper is inference efficiency, we incorporated KANs only as the function approximator for the policy (actor), as this is the component that must be deployed in practice. The value (critic) function remained an MLP for all experiments. We evaluated four training scenarios:
  1. (1)
    MLP actor + MLP critic (both FP),
  2. (2)
    Quantized MLP actor (8-bit) + MLP FP critic,
  3. (3)
    KAN FP actor + MLP FP critic,
  4. (4)
    Quantized KAN actor (8-bit) + MLP FP critic.
To ensure robustness against variance in random initialization, each scenario was trained with 5 different random seeds, for 1 million environment steps per seed. The actor networks were chosen such that the MLP actor had roughly five times more trainable parameters than the KAN actor (Table 6), highlighting that KAN can be more effective even with significantly fewer parameters.
Table 6. Network architectures of the actor and critic.
Model Dimensions Trainable Parameters
MLP Actor [17, 64, 64, 6] 5383
MLP Critic [17, 64, 64, 6] 5383
KAN Actor [17, 6] 1020
5.7.2. Training results
Figure 7 shows the learning curves across the four scenarios. The quantized KAN actor achieves an average return of 2762.2, outperforming both the larger MLP FP and quantized actor baselines (1676.4 and 1558.8) and the full-precision KAN actor (2338.9). These results demonstrate that KANs not only remain robust under aggressive 8-bit quantization, but can also benefit from it, potentially due to regularization effects by fixed bit operations.
5.7.3. Hardware Performance
To evaluate the efficiency of KANs in practical deployment, we compare the hardware cost of the 8-bit quantized KAN actor against the 8-bit quantized MLP actor. The KAN actor was deployed using the KANELÉ framework, while the MLP actor was implemented with hls4ml (Fahim et al., 2021) using the Resource strategy for a fair baseline. We planned to have both models synthesized on a Xilinx xczu7ev-ffvc1156-2-e FPGA in out-of-context mode. However, the 8-bit MLP design exceeds the available FPGA resources, so its performance results are based on HLS estimates rather than the actual implementation. Table 7 summarizes the results. As can be seen, the 8-bit KAN actor achieves significantly lower resource utilization, latency, and power compared to the 8-bit MLP actor, underscoring the advantages of KANs for real-time reinforcement learning control tasks.
In conclusion, the HalfCheetah task is a simulated environment that captures core principles of real-world control tasks and serves as a strong proxy for domains where real-time, resource-efficient policies are essential. These results therefore highlight the suitability of KANs for deployment in settings where such constraints matter (e.g., robotics, embedded control, trading, or quantum computing). Thus, the benchmark serves as a proof-of-concept: KANs can provide competitive or superior RL performance while being dramatically more efficient in terms of size and quantization tolerance.
[…]
Table 7. FPGA performance of the KAN 8-bit model on the HalfCheetah RL task targeting the xczu7ev-ffvc1156-2-e FPGA. The MLP 8-bit actor does not fit on the FPGA, so its power estimate is unavailable and resource usage is reported from HLS estimates, while KAN 8-bit results are obtained after place-and-route in out-of-context mode.
Metric KAN 8-bit MLP 8-bit hls4ml
1M Episode Reward 2762.2 1558.8
Max. Frequency ($F_{\max}$) 884 MHz 500 MHz
Latency 4.5 ns 893 ns
BRAM 0 0
DSP 0 14,346
Flip-Flops (FF) 2,828 460,800
Look-Up Tables (LUT) 1,136 230,400
[…]
Dynamic Power 0.224 nJ/sample $\gg 0.224$ nJ/sample
6. Conclusion and Future Works
We present KANELÉ, a hardware–software co-design framework that maps Kolmogorov–Arnold Networks (KANs) onto a LUT-native computational architecture for FPGAs. Unlike most existing ML hardware–software co-design approaches, KANs are built entirely from learnable 1D activation functions defined on a fixed domain. Each learned activation function $\phi(x)$ is thus not merely approximated by a L-LUT: it is a lookup table. Moreover, the additive structure of KANs enables an especially natural form of pruning: each node can be directly removed from the summation without disrupting the remaining computation. This is in stark contrast to conventional LUT-based neural networks, where LUTs are typically chained together as indices into one another, making the removal of even a single LUT practically impossible without breaking the model. Consequently, mapping a KAN to an FPGA is less a process of compilation and more one of direct instantiation. This paradigm shift—from emulating arithmetic to directly configuring logic—is what unlocks the extreme efficiency of KANELÉ, sidestepping the need for DSPs and BRAMs entirely and aligning the algorithm directly with the hardware’s native capabilities.
[…]
Future Works.
Building on our results, we identify several avenues to broaden the scope and impact of KANELÉ:
  • Broader model families. Extending beyond single KANs to ensembles, temporal and convolutional KANs, graph-based KANs, or transformer-style KANs (“KAN-GPT”) on FPGAs.
  • Alternative orthogonal bases. Moving beyond B-splines by exploring Fourier, wavelet, or rational bases for learnable activations. These alternatives may improve approximation power and training dynamics while remaining LUT-compatible.
  • Practical deployment in control tasks. Our reinforcement learning demo shows that 8-bit KAN policies can outperform larger MLPs while sustaining ultra-low FPGA latency, meeting demands of applications such as quantum error correction, plasma stabilization, adaptive optics, and robotics. Future work targets rapid in-field adaptation through hot-swapping edge tables via partial reconfiguration or LUT updates, enabling lightweight online learning with minimal latency.
Realizing the full potential of this paradigm, however, first requires overcoming the perception that KANs are inherently inefficient in hardware. Our work, KANELÉ, directly refutes this view by aligning the activation-centric KAN formulation with the native strengths of FPGA LUT fabrics. This approach transforms KANs into a practical, high-throughput, and power-efficient inference architecture. We believe this hardware-centric perspective not only solves a critical implementation challenge but also solidifies the promising path toward interpretable, low-resource neural networks that scale from embedded controllers to large scientific instruments, naturally benefiting from software–hardware co-design.
Acknowledgements.
The authors thank the anonymous referees, Vladimir Loncar, Marta Andronic and Ryan Kastner for their valuable comments and helpful suggestions. PH, DH, and AG are supported by the NSF-funded A3D3 Institute (NSF-PHY-2117997).